OpenStudy (anonymous):

4. What are the zeros of the function? What are their multiplicities? F(x)=3x^3+15x^2+18x

OpenStudy (anonymous):

factor a \(3x\) out first

OpenStudy (solomonzelman):

I'll help you find the zeros (this function is simpler than other cubic functions) \[F(x)=3x^3+15x^2+18x~~~~~~~~~~~~~~~~~~set~~i t~~=0\] \[0=3x^3+15x^2+18x~~~~~~~~~~~~~~~~~~~~factor~~out~~of~~x\] \[0=(3x^2+15x+18)x\] so one of the zeros is x=0 \[0=(3x^2+15x+18)~~~~⇒~~~3(x^2+5x+6)=0~~~~⇒~~~x^2+5x+6=0\]

OpenStudy (solomonzelman):

\[(x+3)(x+2)=0~~~~So~~~~~x=-2~~~and~~~x=-3\]so you zeros are \(\huge\color{magenta }{\huge {\bbox[5pt,yellow,border:2px solid red]{0,~~~-2,~~~-3}}}\)

OpenStudy (anonymous):

Thanks